P1 You have been asked to assess the expected financial impact of each of the following...

P1
You have been asked to assess the expected financial impact of each of the following proposals to improve the profitability of credit sales made by your company. Each proposal is independent of the other. Answer all questions. Showing your work may earn you partial credit.

Extend trade credit to some customers that previously have been denied credit because they were considered poor risks.   Sales are projected to increase by $200,000 per year if credit is extended to these new customers. Of the new accounts receivable generated, 6% are projected to be uncollectible. Additional collection costs are projected to be 5% of incremental sales, and production and selling costs are projected to be 78% of sales. Your firm expects to pay a total of 30% of its income after expenses in taxes.

  1. Compute the incremental income after taxes that would result from these projections:
  2. Compute the incremental Return on Sales if these new credit customers are accepted:

If the receivable turnover ratio is expected to be 5 to 1 and no other asset buildup is needed to serve the new customers…

  1. Compute the additional investment in Accounts Receivable
  2. Compute the incremental Return on New Investment
  3. If your company requires a 20% Rate of Return on Investment for all proposals, do the numbers suggest that trade credit should be extended to these new customers? Explain.
Answer #1

Answer 1:

Incremental Income after tax = $15,400

Working:

$200,000 Incremental sales Expenses: Production and selling cost (78% of sales) $156,000 $10,000 Additional collection costs

Answer 2:

Incremental Return on Sales = 7.7%

Working:

Incremental Return on Sales = 15400 / 200000 * 100 = 7.7%

Answer 3:

Additional investment in Accounts Receivable = $40,000

Working:

Additional investment in Accounts Receivable = Sales / receivable turnover ratio = 200000 / 5 = $40,000

Answer 4:

Incremental return on new investment = 15400 / 40000 = 38.5%

Answer 5:

Yes,

Trade credit should be extended to these new customers.

Incremental return on new investment is greater than required rate of return of 20%.